OpenStudy (ell):

Help Pwease! :) find the accumulated value of an investment of $1600 at 8% compounded quarterly for 2 years. A=P(1+r/n) ^nt

OpenStudy (johnweldon1993):

Alright...so compounded quarterly means 4 times a year...that would be our 'n' t = 2 P = 1600 r = .08 So \[\large A = 1600(1 + \frac{.08}{4})^{4 \times 2}\]

OpenStudy (ell):

okay so you put it in your calculator just like that? I got 1874.66

OpenStudy (johnweldon1993):

That's what I got too

OpenStudy (ell):

Okay, shanks! Your a big help :) God Bless!

OpenStudy (johnweldon1993):

Anytime :)